# Retraining with Predicted Hard Labels Provably Increases Model Accuracy

### v1

@@ Appendix G Roadmap of Analysis with Exact Minimizer $\bar{\bm{\theta}}_{0}$
$\Big{|}\langle{\mathbf{x}}^{\prime},\bar{\bm{\theta}}_{0}\rangle-\langle{\mathbf{x}}^{\prime},\widehat{\bm{\theta}}_{0}\rangle\Big{|}\leq\Bigg{(}\frac{C\sqrt{\frac{d+\log(2/\delta)}{n}}}{1-C\sqrt{\frac{d+\log(2/\delta)}{n}}}\Bigg{)}\|{\mathbf{x}}^{\prime}\|\|\widehat{\bm{\theta}}_{0}\|,$ (155)
with a probability of at least $1-\delta$. So if we can bound $\|\widehat{\bm{\theta}}_{0}\|$ with high probability, then we can focus on analyzing $\langle{\mathbf{x}}^{\prime},\widehat{\bm{\theta}}_{0}\rangle$ (which we did previously) and slightly adapting the rest of the analysis to account for this extra perturbation. We can indeed obtain sharp high-probability bounds for $\|\widehat{\bm{\theta}}_{0}\|$ by showing that it is norm-subGaussian as defined in Jin et al., (2019); we skip these calculations here as they are lengthy and tedious. Thus, as stated in Section 4.1, the crux of the proofs is in the analysis of $\langle{\mathbf{x}}^{\prime},\widehat{\bm{\theta}}_{0}\rangle$.
Lemma 10.
For any ${\mathbf{z}}\in\mathbb{R}^{d}$ and ${\mathbf{\Delta}}\in\mathbb{R}^{d\times d}$ such that $\|{\mathbf{\Delta}}\|<1$, we have:
$\|({\mathbf{I}}_{d}+{\mathbf{\Delta}})^{-1}{\mathbf{z}}-{\mathbf{z}}\|\leq\Bigg{(}\frac{\|{\mathbf{\Delta}}\|}{1-\|{\mathbf{\Delta}}\|}\Bigg{)}\|{\mathbf{z}}\|.$ (156)
Proof.
We have:
$\|({\mathbf{I}}_{d}+{\mathbf{\Delta}})^{-1}{\mathbf{z}}-{\mathbf{z}}\|$ $=\|({\mathbf{I}}_{d}+{\mathbf{\Delta}})^{-1}{\mathbf{z}}-({\mathbf{I}}_{d}+{\mathbf{\Delta}})({\mathbf{I}}_{d}+{\mathbf{\Delta}})^{-1}{\mathbf{z}}\|$ (157)
$=\Big{\|}\Big{(}{\mathbf{I}}_{d}-({\mathbf{I}}_{d}+{\mathbf{\Delta}})\Big{)}({\mathbf{I}}_{d}+{\mathbf{\Delta}})^{-1}{\mathbf{z}}\Big{\|}$ (158)
$\leq\|{\mathbf{\Delta}}\|\|({\mathbf{I}}_{d}+{\mathbf{\Delta}})^{-1}{\mathbf{z}}\|.$ (159)
But by the triangle inequality, we also have:
$\|({\mathbf{I}}_{d}+{\mathbf{\Delta}})^{-1}{\mathbf{z}}\|$ $\leq\|({\mathbf{I}}_{d}+{\mathbf{\Delta}})^{-1}{\mathbf{z}}-{\mathbf{z}}\|+\|{\mathbf{z}}\|$ (160)
$\leq\|{\mathbf{\Delta}}\|\|({\mathbf{I}}_{d}+{\mathbf{\Delta}})^{-1}{\mathbf{z}}\|+\|{\mathbf{z}}\|,$ (161)
where eq. 161 follows from eq. 159. Rearranging eq. 161 yields:
$\|({\mathbf{I}}_{d}+{\mathbf{\Delta}})^{-1}{\mathbf{z}}\|\leq\frac{\|{\mathbf{z}}\|}{1-\|{\mathbf{\Delta}}\|}.$ (162)
Plugging this into eq. 159 gives us the desired result. ∎
Appendix H Remaining Experimental Details

### v2

@@ B.2 Lower bound $\widetilde{\alpha}_{0}(\mathbf{x})$
$\mathbb{P}\left(\widetilde{y}_{\ell}^{\prime}\langle\bm{x}_{\ell},{\widehat{\bm{\theta}}}_{0}\rangle<0\right|\mathcal{F}_{\ell})$ $=\Phi\left(-\frac{\widetilde{y}_{\ell}^{\prime}\langle\bm{x}_{\ell},(\sum_{i}\tilde{\xi}_{i}){\bm{{\mu}}}+\bm{\Sigma}^{1/2}\xi_{\ell}y_{\ell}{\bm{z}}_{\ell}\rangle}{\sqrt{n-1}\left\|\bm{\Sigma}^{1/2}\bm{x}_{\ell}\right\|_{\ell_{2}}}\right)$
=$\stackrel{{\scriptstyle(a)}}{{=}}\Phi\left(-\frac{\Big{|}\langle\bm{x}_{\ell},(\sum_{i}\tilde{\xi}_{i}){\bm{{\mu}}}+\bm{\Sigma}^{1/2}\xi_{\ell}y_{\ell}{\bm{z}}_{\ell}\rangle\Big{|}}{\sqrt{n-1}\left\|\bm{\Sigma}^{1/2}\bm{x}_{\ell}\right\|_{\ell_{2}}}\right)$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\frac{1}{2}\exp\left(-\frac{\Big{|}\langle\bm{x}_{\ell},(\sum_{i}\tilde{\xi}_{i}){\bm{{\mu}}}+\bm{\Sigma}^{1/2}\xi_{\ell}y_{\ell}{\bm{z}}_{\ell}\rangle\Big{|}^{2}}{2n\left\|\bm{\Sigma}^{1/2}\bm{x}_{\ell}\right\|_{\ell_{2}}^{2}}\right)$
$\stackrel{{\scriptstyle(c)}}{{\leq}}\frac{1}{2}\exp\left(-\frac{\langle\bm{x}_{\ell},{\bm{{\mu}}}\rangle^{2}(\sum_{i}\tilde{\xi}_{i})^{2}}{4n\left\|\bm{\Sigma}^{1/2}\bm{x}_{\ell}\right\|_{\ell_{2}}^{2}}\right)\exp\left(\frac{\langle\bm{x}_{\ell},\bm{\Sigma}^{1/2}{\bm{z}}_{\ell}\rangle^{2}}{2n\left\|\bm{\Sigma}^{1/2}\bm{x}_{\ell}\right\|_{\ell_{2}}^{2}}\right)$
$\leq\frac{1}{2}\exp\left(-\frac{\langle\bm{x}_{\ell},{\bm{{\mu}}}\rangle^{2}(\sum_{i}\tilde{\xi}_{i})^{2}}{4n\left\|\bm{\Sigma}^{1/2}\bm{x}_{\ell}\right\|_{\ell_{2}}^{2}}\right)\exp\left(\frac{\left\|{\bm{z}}_{\ell}\right\|_{\ell_{2}}^{2}}{2n}\right)$
$\stackrel{{\scriptstyle(d)}}{{\leq}}\frac{1}{2}\exp\left(-\frac{\gamma^{4}(\sum_{i}\tilde{\xi}_{i})^{2}}{4n\lambda_{\max}^{2}\left\|{\bm{z}}_{\ell}\right\|_{\ell_{2}}^{2}}\right)\exp\left(\frac{\left\|{\bm{z}}_{\ell}\right\|_{\ell_{2}}^{2}}{2n}\right)\,.$ (E.13)
Here, $(a)$ follows by the choice of $\widetilde{y}^{\prime}_{\ell}$; $(b)$ holds by using (Abramowitz and Stegun,, 1968, Equation (7.1.13)):
$\Phi^{c}(t)\leq\sqrt{\frac{2}{\pi}}\frac{e^{-t^{2}/2}}{t+\sqrt{t^{2}+\frac{8}{\pi}}}\leq\frac{1}{2}e^{-t^{2}/2}\,,$
for all $t\geq 0$. In addition $(c)$ holds since $(a+b)^{2}\geq\frac{a^{2}}{2}-b^{2}$. Also, $(d)$ holds because under our data model we have $\bm{\Sigma}^{1/2}\bm{x}_{\ell}=\bm{\Sigma}\bm{z}_{\ell}$ and so $\left\|\bm{\Sigma}^{1/2}\bm{x}_{\ell}\right\|_{\ell_{2}}\leq\lambda_{\max}\left\|\bm{z}_{\ell}\right\|_{\ell_{2}}$. Further, $|\langle\bm{x}_{\ell},{\bm{{\mu}}}\rangle|=|y(1+u_{\ell})\left\|{\bm{{\mu}}}\right\|_{\ell_{2}}^{2}|\geq\gamma^{2}$ because $u_{\ell}\geq 0$. Putting these bounds together we have
$\frac{\langle\bm{x}_{\ell},{\bm{{\mu}}}\rangle}{\left\|\bm{\Sigma}^{1/2}\bm{x}_{\ell}\right\|_{\ell_{2}}}\geq\frac{\gamma^{2}}{\lambda_{\max}\left\|{\bm{z}}_{\ell}\right\|_{\ell_{2}}}\,.$
We proceed by taking expectation of the right-hand side of (E.13) with respect to $\tilde{\xi}_{1},\dotsc,\tilde{\xi}_{n}$. Define the shorthand $\lambda:=\frac{\gamma^{4}}{4n\lambda_{\max}^{2}\left\|{\bm{z}}_{\ell}\right\|_{\ell_{2}}^{2}}$. Fix an arbitrary $\lambda_{0}>0$ for now (we will determine its value later) and define truncated parameter $\bar{\lambda}=\min(\lambda,\lambda_{0})$. We have
$\operatorname{\mathbb{E}}\left[e^{-\lambda(\sum_{i}\tilde{\xi}_{i})^{2}}\right]$ $\leq\operatorname{\mathbb{E}}\left[e^{-\bar{\lambda}(\sum_{i}\tilde{\xi}_{i})^{2}}\right]$ (E.14)
$=\operatorname{\mathbb{E}}\left[e^{-\bar{\lambda}\sum_{i}(1+u_{i})^{2}}e^{-\bar{\lambda}\sum_{i\neq j}\xi_{i}\xi_{j}(1+u_{i})(1+u_{j})}\right]$
$\leq e^{-\bar{\lambda}n}\prod_{i\neq j}\operatorname{\mathbb{E}}[e^{-\bar{\lambda}\xi_{i}(1+u_{i})}]\operatorname{\mathbb{E}}[e^{-\bar{\lambda}\xi_{j}(1+u_{j})}]$
$=e^{-\bar{\lambda}n}(\operatorname{\mathbb{E}}[e^{-\bar{\lambda}\xi(1+u)}])^{n(n-1)}$
$=e^{-\bar{\lambda}n}(\operatorname{\mathbb{E}}[(1-p)e^{-\bar{\lambda}(1+u)}+pe^{\bar{\lambda}(1+u)}])^{n(n-1)}$
$\leq e^{-\bar{\lambda}n}\left((1-p)e^{-\bar{\lambda}}+pe^{\bar{\lambda}}\operatorname{\mathbb{E}}[e^{\bar{\lambda}u}]\right)^{n(n-1)}$
$\leq e^{-\bar{\lambda}n}\left((1-p)e^{-\bar{\lambda}}+pe^{\bar{\lambda}}e^{\bar{\lambda}^{2}/2}\right)^{n(n-1)}$
$=e^{-\bar{\lambda}n^{2}}\left(1-p+pe^{2\bar{\lambda}+\bar{\lambda}^{2}/2}\right)^{n(n-1)}\,.$ (E.15)
Here, the first and the second inequality follows from $u_{i}\geq 0$. The third inequality holds since $u$ has unit sub-gaussian norm and so it moment-generating function is bounded as $\operatorname{\mathbb{E}}[e^{\lambda u}]\leq e^{\lambda^{2}/2}$.
We next further upper bound the right-hand side of E.15 to get a simpler expression. In doing that we use the following lemma.
Lemma E.3
For $t\leq 1/2$ we have $e^{t}-1\leq t+t^{2}$.
By choosing $\lambda_{0}\leq 0.2$ we have $22\lambda_{0}+\lambda_{0}^{2}/2\leq 1/$ 2 and so $22\bar{\lambda}+\bar{\lambda}^{2}/2\leq 1/$ 2. By virtue of the above lemma we have
$e^{2\bar{\lambda}+\bar{\lambda}^{2}/2}-1$ $\leq 2\bar{\lambda}+\bar{\lambda}^{2}/2+\left(2\bar{\lambda}+\bar{\lambda}^{2}/2\right)^{2}$
$=2\bar{\lambda}+\bar{\lambda}^{2}/2+4\bar{\lambda}^{2}+\bar{\lambda}^{4}/4+2\bar{\lambda}^{3}$
$\leq 2\bar{\lambda}+\bar{\lambda}\left(\lambda_{0}/2+4\lambda_{0}+\lambda_{0}^{3}/4+2\lambda_{0}^{2}\right)$
$\leq(2+6\lambda_{0})\bar{\lambda}\,,$
where we used $\lambda_{0}\leq 0.2$ in the last step. Next using the inequality $1+x\leq e^{x}$ for $x\geq 0$, we get
$(e^{2\bar{\lambda}+\bar{\lambda}^{2}/2}-1)p+1\leq(2+6\lambda_{0})p\bar{\lambda}+1\leq e^{(2+6\lambda_{0})\bar{\lambda}p}\,.$
By using this bound in (E.15) we obtain
$\operatorname{\mathbb{E}}\left[e^{-\lambda(\sum_{i}\tilde{\xi}_{i})^{2}}\right]\leq e^{-\bar{\lambda}n^{2}}e^{(2+6\lambda_{0})\bar{\lambda}pn^{2}}=e^{-(1-2p^{\prime})\bar{\lambda}n^{2}}\,,$ (E.16)
with $p^{\prime}:=(1+3\lambda_{0})p$.
Using the above in (E.13) and then taking expectation with respect to ${\bm{z}}_{\ell}$ we get
$\mathbb{P}\left(\widetilde{y}_{\ell}^{\prime}\langle\bm{x}_{\ell},{\widehat{\bm{\theta}}}_{0}\rangle<0\right)$ $\leq\frac{1}{2}\operatorname{\mathbb{E}}\left[\exp\left(-(1-2p^{\prime})\bar{\lambda}n^{2}\right)\exp\left(\frac{\left\|{\bm{z}}_{\ell}\right\|_{\ell_{2}}^{2}}{2n}\right)\right]$
$\leq\frac{1}{2}\operatorname{\mathbb{E}}\left[\exp\left(-2(1-2p^{\prime})\bar{\lambda}n^{2}\right)\right]^{1/2}\operatorname{\mathbb{E}}\left[\exp\left(\frac{\left\|{\bm{z}}_{\ell}\right\|_{\ell_{2}}^{2}}{n}\right)\right]^{1/2}\,,$ (E.17)
by applying the Cauchy–Schwarz inequality.
Using the moment generating function of $\chi^{2}_{d}$ distribution, we have
$\operatorname{\mathbb{E}}\left[\exp\left(\frac{\left\|{\bm{z}}_{\ell}\right\|_{\ell_{2}}^{2}}{n}\right)\right]=\left(1-\frac{2}{n}\right)^{-d/2}\leq e^{2d/n}\,,$ (E.18)
for $n\geq 4$. Here, we use the inequality $(1-x)^{-1}\leq e^{2x}$ for $x\in[0,1/2]$.
We continue by bounding the first term on the right-hand side of (E.17). Define the event $\mathcal{E}_{1}$ as follows:
$\mathcal{E}_{1}:=\left\{\left\|{\bm{z}}_{\ell}\right\|_{\ell_{2}}\leq\sqrt{2d}\right\}\,.$ (E.19)
We then have
$\operatorname{\mathbb{E}}\left[\exp\left(-2(1-2p^{\prime})\bar{\lambda}n^{2}\right)\right]$ $=\operatorname{\mathbb{E}}\left[\exp\left(-2(1-2p^{\prime})\bar{\lambda}n^{2}\right)(\mathbb{1}_{\mathcal{E}_{1}}+\mathbb{1}_{\mathcal{E}_{1}^{c}})\right]$
$\leq\operatorname{\mathbb{E}}\left[\exp\left(-2(1-2p^{\prime})\bar{\lambda}n^{2}\right)\mathbb{1}_{\mathcal{E}_{1}}\right]+\operatorname{\mathbb{E}}[\mathbb{1}_{\mathcal{E}_{1}^{c}}]$
=$\stackrel{{\scriptstyle(a)}}{{=}}\operatorname{\mathbb{E}}\left[\exp\left(-\frac{\gamma^{4}}{4\lambda_{\max}^{2}d}(1-2p^{\prime})n\right)\mathbb{1}_{\mathcal{E}_{1}}\right]+\mathbb{P}(\mathcal{E}_{1}^{c})$
$\leq\exp\left(-\frac{\gamma^{4}}{4\lambda_{\max}^{2}}(1-2p^{\prime})\frac{n}{d}\right)+\mathbb{P}(\mathcal{E}_{1}^{c})$ (E.20)
Note that in step $(a)$, we used the fact that on the event $\mathcal{E}_{1}$, we have $\lambda\geq{\gamma^{4}}/(8\lambda_{\max}^{2}nd)$ and we choose $\lambda_{0}={\gamma^{4}}/(8\lambda_{\max}^{2}nd)$. This way, we have $\bar{\lambda}=\min(\lambda,\lambda_{0})={\gamma^{4}}/(8\lambda_{\max}^{2}nd)$. Also note that by our assumption in the statement of the lemma, we have $\lambda_{0}\leq 0.2$ which is the condition assumed in deriving (E.16).
We next bound $\mathbb{P}(\mathcal{E}_{1}^{c})$. Since ${\bm{z}}_{\ell}\sim{\mathcal{N}}(0,\bm{I}_{d})$, $\left\|\bm{z}_{\ell}\right\|_{\ell_{2}}^{2}$ is $\chi^{2}$ distribution with $d$-degrees of freedom. Using the tail bound of $\chi^{2}$ distribution (see e.g, (Ghosh,, 2021, Theorem 3)) we have $\mathbb{P}(\left\|{\bm{z}}_{\ell}\right\|_{\ell_{2}}^{2}\geq 2d)\leq e^{-d/8}$ and so
$\mathbb{P}(\mathcal{E}_{1}^{c})\leq\mathbb{P}(\left\|{\bm{z}}_{\ell}\right\|_{\ell_{2}}^{2}\geq 2d)\leq e^{-d/8}.$
Using the above in (E.20) we obtain
$\operatorname{\mathbb{E}}\left[\exp\left(-2(1-2p^{\prime})\bar{\lambda}n^{2}\right)\right]$ $\leq\exp\left(-\frac{\gamma^{4}}{4\lambda_{\max}^{2}}(1-2p^{\prime})\frac{n}{d}\right)+e^{-d/8}.$ (E.21)
By combining (E.21) and (E.18) with (E.17) we arrive at
$\mathbb{P}\left(\widetilde{y}_{\ell}^{\prime}\langle\bm{x}_{\ell},{\widehat{\bm{\theta}}}_{0}\rangle<0\right)$ $\leq\frac{1}{2}\left(\exp\left(-\frac{\gamma^{4}}{4\lambda_{\max}^{2}}(1-2p^{\prime})\frac{n}{d}\right)+e^{-d/8}\right)^{1/2}e^{d/n}$
$\leq\frac{1}{2}\left(\exp\left(-\frac{\gamma^{4}}{8\lambda_{\max}^{2}}(1-2p^{\prime})\frac{n}{d}\right)+e^{-d/16}\right)e^{d/n}\,,$
which completes the proof of lemma.
E.2 Proof of Lemma E.2
Under our data model we have $\bm{x}_{\ell}=y_{\ell}(1+u_{\ell}){\bm{{\mu}}}+\bm{\Sigma}^{1/2}{\bm{z}}_{\ell}$. Substituting for $\bm{x}_{\ell}$ in the expression of $T_{\ell}$, we have
$T_{\ell}$ $={\rm sign}\left(\langle y_{\ell}(1+u_{\ell}){\bm{{\mu}}}+\bm{\Sigma}^{1/2}{\bm{z}}_{\ell},\beta{\bm{{\mu}}}+\xi_{\ell}y_{\ell}\bm{\Sigma}^{1/2}{\bm{z}}_{\ell}\rangle\right)\langle y\bm{x},y_{\ell}(1+u_{\ell}){\bm{{\mu}}}+\bm{\Sigma}^{1/2}{\bm{z}}_{\ell}\rangle$
$={\rm sign}\left(y_{\ell}(1+u_{\ell})\beta\gamma^{2}+\xi_{\ell}y_{\ell}\left\|\bm{\Sigma}^{1/2}{\bm{z}}_{\ell}\right\|_{\ell_{2}}^{2}\right)\langle y\bm{x},y_{\ell}(1+u_{\ell}){\bm{{\mu}}}+\bm{\Sigma}^{1/2}{\bm{z}}_{\ell}\rangle\,,$
using the fact that $\bm{\Sigma}^{1/2}{\bm{{\mu}}}=0$. Taking expectation we get
$\operatorname{\mathbb{E}}[T_{\ell}|\beta]=\operatorname{\mathbb{E}}\left[{\rm sign}\left(y_{\ell}(1+u_{\ell})\beta\gamma^{2}+\xi_{\ell}y_{\ell}\left\|\bm{\Sigma}^{1/2}{\bm{z}}_{\ell}\right\|_{\ell_{2}}^{2}\right)\langle y\bm{x},y_{\ell}(1+u_{\ell}){\bm{{\mu}}}\rangle\;\Big{|}\beta\right]$
since the other term will be an odd function of ${\bm{z}}_{\ell}$. Letting
$q_{0}:=\mathbb{P}\left((1+u_{\ell})\beta\gamma^{2}+\xi_{\ell}\left\|\bm{\Sigma}^{1/2}{\bm{z}}_{\ell}\right\|_{\ell_{2}}^{2}<0\right)\,,$
we get
$\operatorname{\mathbb{E}}[T_{\ell}|\beta]=(1-2q_{0})\langle y\bm{x},(1+u_{\ell}){\bm{{\mu}}}\rangle\geq(1-2q_{0})\langle y\bm{x},{\bm{{\mu}}}\rangle\,,$
because $u_{\ell}>0$ and $\langle y\bm{x},{\bm{{\mu}}}\rangle>0$, given the positive margin in our data model, and if $q_{0}\leq\frac{1}{2}$. So what remains is to show that $q_{0}\leq q\leq 1/2$.
We write
$q_{0}$ $\leq\mathbb{P}\left((1+u_{\ell})\beta\gamma^{2}-\lambda_{\max}\left\|{\bm{z}}_{\ell}\right\|_{\ell_{2}}^{2}<0\right)$
$\leq\mathbb{P}\left(\frac{\beta\gamma^{2}}{\lambda_{\max}}<\left\|{\bm{z}}_{\ell}\right\|_{\ell_{2}}^{2}\right)\quad\quad\quad\quad\quad\quad\quad\quad\text{(since $u_{\ell}>0$)}$
$<\exp\left(-\frac{\beta\gamma^{2}}{20\lambda_{\max}}\right)\,,$
if $\frac{\beta\gamma^{2}}{2d\lambda_{\max}}>1$, where the last step follows from the observation that $\left\|{\bm{z}}_{\ell}\right\|_{\ell_{2}}^{2}\sim\chi^{2}_{d}$ and using the tail bound of $\chi^{2}_{d}$(Ghosh,, 2021, Theorem 3)). So by definition of $q_{0}$ we showed that
$q_{0}\leq q:=\exp\left(-\frac{\beta\gamma^{2}}{20\lambda_{\max}}\right)\leq e^{-d/10}<\frac{1}{2}\,,$
which completes the proof of lemma.
Appendix F Proof of Theorem 4.9
Using Theorem 4.8, we have ${\sf acc}({\widehat{\bm{\theta}}}_{1})\geq\operatorname{\mathbb{E}}[\alpha_{1}({\bm{x}})]$. Based on our data model, $\bm{x}=y(1+u){\bm{{\mu}}}+\bm{\Sigma}^{1/2}{\bm{z}}$. Consider the probabilistic event $\mathcal{E}_{2}:=\{{\bm{z}}:\left\|{\bm{z}}\right\|_{\ell_{2}}^{2}\leq 2d\}$. For every $\eta>0$, we have
$\operatorname{\mathbb{E}}\left[\exp\left(-\frac{\eta\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{\langle\bm{x},{\bm{{\mu}}}\rangle^{2}+\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\right)\right]$ $=\operatorname{\mathbb{E}}\left[\exp\left(-\frac{\eta\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{\langle\bm{x},{\bm{{\mu}}}\rangle^{2}+\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\right)(\mathbb{1}_{\mathcal{E}_{2}}+\mathbb{1}_{\mathcal{E}_{2}^{c}})\right]$
$\leq\exp\left(-\frac{\eta\gamma^{4}}{\gamma^{4}+2\lambda_{\max}^{2}d}\right)+\mathbb{P}(\mathcal{E}_{2}^{c})\,,$
where the inequality holds because $\langle\bm{x},{\bm{{\mu}}}\rangle=y(1+u)\gamma^{2}$ and so $\langle\bm{x},{\bm{{\mu}}}\rangle^{2}\geq\gamma^{4}$ (since $u$ is a non-negative random variable). In addition,
$\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}=\left\|\bm{\Sigma}{\bm{z}}\right\|_{\ell_{2}}^{2}\leq\lambda_{\max}^{2}\left\|{\bm{z}}\right\|_{\ell_{2}}^{2}\leq 2\lambda_{\max}^{2}d$
under the event $\mathcal{E}_{2}$. Further, $\left\|{\bm{z}}\right\|_{\ell_{2}}^{2}\sim\chi^{2}_{d}$ and so using tail bounds for $\chi^{2}_{d}$, see e.g., (Ghosh,, 2021, Theorem 3), we have $\mathbb{P}(\mathcal{E}_{2}^{c})\leq e^{-d/8}$. Putting things together, we obtain
$\operatorname{\mathbb{E}}\left[\exp\left(-\frac{\eta\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{\langle\bm{x},{\bm{{\mu}}}\rangle^{2}+\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\right)\right]\leq\exp\left(-\frac{\eta\gamma^{4}}{\gamma^{4}+2\lambda_{\max}^{2}d}\right)+e^{-d/8}\,.$
The claim of theorem follows by setting $\eta=\frac{n}{64}(1-2q^{\prime})^{2}$.
Appendix G Remaining Experimental Details